Fermi National Laboratory


Accelerator Update

Friday April 30
The day shift began with Operations monitoring store 3458 and the Recycler stash, with the Antiproton Source (Pbar) stacking, and with MiniBooNE and the Meson MTest experiments taking beam.

At 9:10 AM, an expert added water to the Main Injector's (MI) RF cavity LCW system. At 9:18 AM the MI RF system tripped off due to a LCW pressure alarm. Operations reported that the beam was back and the LCW pressure had normalized by 9:27 AM.

An MTest expert reported at 9:58 AM that the experiment was ready to receive fast spill. Operations established fast spill beam to MTest at 10:47 AM. An expert tuned the beam for another hour or so.

Linac experts conducted a 750 KeV Line chopper timing study at 10:56 AM. The study lasted about forty minutes.

A Linac Klystron RF station (KRF5) tripped off at 12:09 PM and then again five minutes later. A Linac expert investigated and discovered that another Linac expert was tuning the transition buncher phase. Linac resumed sending beam at 12:43 PM.

At 1:30 PM, Pbar suffered from an emittance blowup. This halted stacking for about ten minutes while an operator tuned.

All but one of the MI RF stations tripped off at 3:53 PM due to a phase stability problem.

Operations reported at 6 PM that they had resolved another Pbar emittance blow up.

A Booster RF device tripped off at 7:24 PM due to a bad power supply. Operators replaced the supply.

The MTest experiments requested that fast spill beam be disabled at 7:29 PM. Operations enabled the slow spill beam at 8:45 PM.

Operators made a small Tevatron (TeV) collimator move 7:36 PM that improved the TeV beam.

A Pbar expert began a reverse protons study at 8:54 PM.

Saturday May 1
The midnight shift began with Operations monitoring store 3458 and a stash, with Pbar conducting studies, and with MiniBooNE and MTest taking beam.

At 4:35 AM, Operations reported that beam wasn't getting into the Accumulator. Pbar resumed stacking at 5:01 AM after operators resolved a data error.

At 7:24 AM, Pbar suffered from an interlock timing problem. Operators adjusted events in the Time Line Generator and stacking resumed.

MTest experimenter requested fast spill beam. Operations established fast spill at 10:12 AM.

At 11:02 AM, Operations terminated store 3458 and then put the TeV into shot setup at 11:58 AM. Switchyard beam was disabled.

Operations established store 3460 at 1:46 PM with an initial luminosity of 57.2E30. MTest resumed taking beam.

Operations attempted to have Pbar resume stacking at 2:39 PM, but the stacking rate was poor. Operation contacted an expert who had them make a phase change, which resolved the problem at 3:32 PM.

Operations put the Recycler into shot setup to stash antiprotons from Pbar at 6:06 PM. The transfers ended and stacking resumed at 6:49 PM.

Sunday May 2
The midnight shift began with Operation monitoring store 3460 and the stash, with Pbar stacking, and with MiniBooNE and MTest taking beam.

MTest finished using fast spill at 12:04 AM and operators reestablished slow spill to both MTest and MCenter experiments.

At 4:25 AM, a MI device (MI60N) suffered an overcurrent trip. It took about twelve minutes before Operations could reset this device. At 5:07 AM, it tripped off again. Operators called in experts to investigate and they found a bad SCR and a water leak that caused three firing circuit cards to go bad. They completed their repairs by 7:46 AM.

A Linac vacuum pump tripped off at 10:36 AM. It reset okay.

A Linac RF station (KRF5) tripped off at 11:11 AM.

A Linac RF station (KRF4) tripped off at 2:54 PM.

At 6:33 PM, Operations reported that the Pbar stacking rate had been getting worse throughout the shift. Operators have been investigating and making small gains, but hadn't found any reason for the problems. Experts were notified.

MTest experimenters said they were finished for the night and requested their beam be disabled at 7:31 PM.

Experimenters from the Collider Detector at Fermilab (CDF) reported at 7:38 PM that their silicon detector went down. They were consulting with experts.

Operations terminated store 3460 at 8:30 PM and then put the TeV into shot setup.

Operations put the TeV through a dry squeeze at 8:48 PM, but suffered a separator spark at 8:51 PM. Operations then conducted several more dry squeezes for a spark investigation.

Operations reported at 10:34 PM a controls problem (CFSS) that was locking up readbacks and sequencers. Operators investigated.

Monday May 3
The midnight shift began with the TeV in shot setup, with Operations monitoring the stash, with operators trying to resolve CFSS problems, with MiniBooNE taking beam, but with MCenter off due to shot setup.

At 12:19 AM, operators consulted with a controls expert.

Operations established store 3464 at 3:47 AM with an initial luminosity of 52.30E30. Pbar resumed stacking.

Operations resumed sending beam to MCenter at 5:28 AM after re-interlocking the MC7 enclosure.

Operations prepared Recycler for stashing antiprotons at 7:20 AM.

Plans for Today
The plan for to day is to stack and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 5/3/2004   email Fermilab